OpenStudy (fanduekisses):

Find an equation in standard form for the hyperbola with vertices at (0, ±6) and asymptotes at y = ±3/2x

OpenStudy (anonymous):

any idea?

OpenStudy (fanduekisses):

vertical right?

OpenStudy (anonymous):

don't forget asymptotes y = \(\pm\)b/a

OpenStudy (fanduekisses):

ohh so it is y^2/4 - x^2/9 ?

OpenStudy (anonymous):

not yet!! you still have a step to figure out which is the form from 

OpenStudy (fanduekisses):

vertical

OpenStudy (anonymous):

vertex (0, 6) means the vertical form, right!!

OpenStudy (anonymous):

ye, you get the correct answer

OpenStudy (fanduekisses):

yay thanks ^_^

OpenStudy (anonymous):

oh, we go backward, not that I am sorry

OpenStudy (anonymous):

the vertical form have the asymptote formula is y =\(\pm a/b\) so, a = 3, b =2 the standard form is \[\dfrac{y^2}{9}-\dfrac{x^2}{4}=1\]

OpenStudy (anonymous):

*\(y=\pm \dfrac{a}{b}x\) I am so sorry for my restlessness.